When silver sulfate (Ag2SO4) is dissolved in water at 25 °C, until saturation is reached, the Ag+ cation equilibrium concentration is 3.0 x 10 ^-5 M.
a)
Determine the value of the solubility constant for Ag2SO4
(b) Determine the maximum amount (in grams) of AgNO3, a completely water-soluble salt, which can be dissolved in 3.2 L of a H2SO4 solution without precipitating Ag2SO4

Note: Actually the molarity of H2SO4 is not given in the exercise. Is it really necessary for the exercise resolution? If it is needed can you please pick a random molarity that you think fits the exercise. At least that way I can learn how to solve this kind of exercises. Thanks in advance!
Answer #1

Part (a).

Ag2SO4 <----> 2Ag+(aq) + SO42-(aq)

Here, [Ag+] = 3*10-5 M

Now, the solubility product constant (Ksp) of Ag2SO4 = [Ag+]2[SO42-] = (2s)2*(s)

Where, s = molar solubility of Ag2SO4 = 3*10-5 M/2 = 1.5*10-5 M

Hence, Ksp = (3*10-5)2*(1.5*10-5) = 1.35*10-14

Part (b).

Up to the value of [Ag+] = 3*10-5 M, the precipitation Ag2SO4 will not take place.

i.e. The moles of Ag+ = 3*10-5 mol/L * 3.2 L = 9.6*10-5 mol

Therefore, the required mass of AgNO3 = 9.6*10-5 mol * 169.87 g/mol = 0.0163 g